# SWI: Speaking with Intent in Large Language Models

### v1

@@ 4.3 Experimental Comparison
CoT aims to elicit LLM reasoning using the answer-trigger prompt $\Phi_{i}^{\text{CoT}}$ as “Let’s think step by step”, while ARR, with a more systematic design, utilizes an enhanced $\Phi_{i}^{\text{ARR}}$ as “Let’s analyze the intent of the question, find relevant information, and answer the question with step-by-step reasoning.” Finally, PS applies the following prompt $\Phi_{i}^{\text{PS}}$ to construct plans before problem-solving: “Let’s first understand the problem and devise a plan to solve the problem. Then, let’s carry out the plan and solve the problem step by step.” With answer-trigger prompts $\Phi_{i}$, the generation process (Eq. 1) is given by
@@
4.4 Experimental Results
SWI vs. Baseline.
As shown in Table 2, SWI improves Baseline (generation without intent) performance by an average of +3.34 points, showing its effectiveness in enhancing LLM’s problem analysis and mathematical reasoning abilities.
Method GSM8K GSM8K-P MATH500 Competition Avg.
Baseline 78.47 81.14 36.80 10.56 51.74
SWI (ours) 81.20 84.20 41.60 13.33 55.08
Table 2: Mathematical Reasoning Results. The exact matching scores (%) of LLaMA-3-8B-Instruct on multiple mathematical reasoning datasets with or without Speaking with Intent.
SWI vs. Answer-Trigger Prompting Methods.
Table 3 presents the results of our SWI method and three answer-trigger prompting methods: CoT, PS, and ARR. We observe that SWI significantly outperforms PS, suggesting that the dynamic generation of intents in SWI provides more effective analytical guidance than the static construction of multiple plans at the outset, as in PS. Furthermore, compared to the reasoning-eliciting methods CoT and ARR, SWI demonstrates superior performance over CoT and achieves competitive results with ARR. Notably, SWI surpasses ARR in solving competition-level mathematics problems, which are more challenging and thus indicate the potential of SWI.
Method GSM8K GSM8K-P MATH500 Competition Avg.
CoT (Kojima et al., 2022) 79.30 82.22 42.60 7.78 52.98
PS (Wang et al., 2023) 73.54 76.43 39.20 8.61 49.45
ARR (Yin & Carenini, 2025) 81.96 85.36 44.40 10.28 55.50
SWI (ours) 81.20 84.20 41.60 13.33 55.08
Table 3: Mathematical Reasoning Results. The exact matching scores (%) of LLaMA-3-8B-Instruct on multiple mathematical reasoning datasets using our SWI method or answer-trigger prompting methods. The underlined values indicate the second-best scores.
5 Generalizability of Speaking with Intent
The main experiments have verified the effectiveness of Speaking with Intent, this section further studies the generalizability of the proposed SWI method on different categories of tasks: multiple-choice QA and text summarization.
5.1 Multiple-choice Question Answering
Many challenging benchmarks are designed as multiple-choice QA tasks, where the model is asked to select the most appropriate one from the given options to answer the question.
5.1.1 QA Datasets
Our SWI method is evaluated on various QA datasets involving reading comprehension (Liu et al., 2020), commonsense reasoning (Talmor et al., 2019; Sap et al., 2019), world knowledge (Mihaylov et al., 2018), and reasoning-intensive multitask understanding tasks with multiple subtasks covering diverse topics (Clark et al., 2018; Suzgun et al., 2023; Hendrycks et al., 2021a; Wang et al., 2024e). The dataset statistics are presented in Table 4.
[…]
Reading LogiQA (Liu et al., 2020) Test 651 185
Commonsense CSQA (Talmor et al., 2019) Valid 1,221 127
SIQA (Sap et al., 2019) Valid 1,954 117
World Knowledge OBQA (Mihaylov et al., 2018) Test 500 127
ARC (Clark et al., 2018) Test 3,548 144
Multitask Understanding BBH (Suzgun et al., 2023) Test 5,511 203
MMLU (Hendrycks et al., 2021a) Test 13,842 193
MMLU-Pro (Wang et al., 2024e) Test 12,032 270
[…]
5.1.2 QA Results
Table 5 shows that our SWI method consistently improves the Baseline (generation without intent) by a large margin, i.e., +10.08 points on average. The results demonstrate the efficacy of Speaking with Intent in reasoning-intensive question answering.
Method Reading Commonsense World Knowledge Multitask Understanding Avg.
LogiQA CSQA SIQA OBQA ARC BBH MMLU MMLU-Pro
Baseline 20.74 66.91 29.58 70.60 76.44 56.65 52.40 39.27 51.57
SWI (ours) 33.03 72.73 46.37 81.60 87.82 64.29 62.19 45.16 61.65
Table 5: Multiple-choice QA Results. The accuracy scores (%) of LLaMA3-8B-Instruct on various multiple-choice QA datasets with or without Speaking with Intent.
@@ 5.2 Text Summarization
The results on math and QA benchmarks have solidified the efficacy of SWI on reasoning-intense tasks. Beyond reasoning tasks, we hypothesize that Speaking with Intent benefits natural language generation tasks like summarization by more explicitly analyzing the source document point by point and better planning the generation of the final summary.
5.2.1 Summarization Datasets
To verify the hypothesis, we test the effect of SWI on various text summarization datasets. CNN/DailyMail (CDM) (Hermann et al., 2015; See et al., 2017), Extreme summarization (XSum) (Narayan et al., 2018), and XL-Sum (Hasan et al., 2021) contain diverse article-summary pairs from news articles. SAMSum (Gliwa et al., 2019) and DialogSum (Chen et al., 2021) are dialogue summarization datasets. WikiLingua (Ladhak et al., 2020) extracts article and summary pairs from wikiHow. The dataset statistics are presented in Table 6.
[…]
News Article CDM (Hermann et al., 2015) Test 11,490 920
XSum (Narayan et al., 2018) Test 11,334 542
XL-Sum (Hasan et al., 2021) Test 11,535 613
Dialogue SAMSum (Gliwa et al., 2019) Test 819 209
DialogSum (Chen et al., 2021) Test 1,500 263
Wiki Article WikiLingua (Ladhak et al., 2020) Test 3,000 525
[…]
5.2.2 Summarization Results
Automatic Evaluation of Summaries.
@@ 5.2.2 Summarization Results
We evaluate the quality of summaries using the ROUGE score (Lin, 2004), which counts the overlaps of the generated summaries and reference summaries. Specifically, we average the ROUGE-1 (unigrams), ROUGE-2 (bigrams), ROUGE-L (longest common subsequences), and ROUGE-LSum (sentence-level ROUGE-L) scores as the final ROUGE score. As shown in Table 7, our SWI method consistently surpasses the Baseline on ROUGE scores, confirming its effectiveness in enhancing the quality of text summaries.
Method News Article Dialogue Wiki Article Avg.
CDM XSum XL-Sum SAMSum DialogSum WikiLingua
Baseline 23.38 11.90 11.29 24.14 16.92 15.01 17.11
SWI (ours) 24.00 13.82 13.50 24.37 20.64 17.86 19.03
Table 7: Text Summarization Results. The ROUGE scores (%) of the LLaMA3-8B-Instruct model on various text summarization datasets with or without Speaking with Intent (SWI).
5.2.3 Fact Checking of Summaries
LLMs frequently generate hallucinated content (Ji et al., 2023; Li et al., 2024a), which can not be detected by lexical metrics like ROUGE. To assess this issue, we adopt a more semantically sophisticated fact-checking metric proposed by Hwang et al. (2025), which quantifies factual consistency by calibrating the extent of fabricated statements (low precision) and omitted factual information (low recall). Specifically, we use GPT-4o-mini (Hurst et al., 2024) to decompose both generated and reference summaries into atomic fact sets and measure their coverage to quantify factual consistency.
We evaluate $100$ samples from each summarization dataset using this fact-checking metric, with results presented in Table 8. Speaking with Intent (SWI) generates more concise and accurate summaries, improving precision, whereas baseline summaries tend to be more lengthy and verbose, resulting in higher recall scores. Overall, SWI consistently outperforms the Baseline generation in terms of F1 score.
Dataset Method Precision Recall F1 Dataset Method Precision Recall F1
CDM Baseline 26.06 76.28 36.37 SAMSum Baseline 39.95 75.92 46.90
SWI 34.22 55.89 37.79 SWI 47.82 69.88 50.88
XSum Baseline 11.06 48.38 15.15 DialogSum Baseline 23.99 57.08 29.55
SWI 14.77 37.30 16.29 SWI 34.92 45.19 31.20
XL-Sum Baseline 8.96 61.88 13.79 WikiLingua Baseline 23.33 65.55 30.63
SWI 12.96 46.72 16.51 SWI 32.40 54.98 35.78
@@ 5.2.3 Fact Checking of Summaries
Table 8: Summarization evaluation results. We compare the atomic facts drawn from the generated and reference summaries, and compute recall, precision, and F1 scores (%).
5.2.4 Case Study
To provide more insights into the benefits of Speaking with Intent, we conduct a case study on the CDM summarization dataset (Hermann et al., 2015; See et al., 2017). Figure 3 showcases the SWI outputs, where the generated intent is well formulated and articulated, guiding a point-by-point summarization process that leads to a final summary that is accurate, concise, and abstractive, while effectively capturing the key information of the source article. This probably leads to the high Precision and F1 scores of SWI in Table 8, suggesting the validity of Speaking with Intent in text generation tasks.
@@ 5.2.4 Case Study
Figure 3: Case Study. Task input and SWI output on text summarization task.
6 Intent Quality Evaluation
@@
Although we have shown that Speaking with Intent boosts performance across a broad range of tasks, verifying the quality of generated intents is also significant. Thus, we hire human evaluators to assess the quality of generated intent across three criteria: coherence, effectiveness, and interpretability. Coherence measures how well the intent guides analysis and reasoning, effectiveness evaluates its contribution to problem-solving, and interpretability assesses its role in enhancing user understanding of the generated content. [^3] Please refer to Appendix C for more details on human evaluation. For each instance, human evaluators are provided with evaluation instructions, task input (e.g., the math problem, question with options, or source article), SWI-generated output, and assessment check boxes. They are then asked to evaluate the following aspects:
@@
  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for solving the problem?
@@
Evaluation scores range from 1 (Bad), 2 (Fair), to 3 (Good). Agreement ratios are calculated as follows: 1 if all three evaluators agree, 0.5 if two agree, and 0 if all scores differ. As shown in Table 9, human evaluation scores for all aspects across datasets exceed 2, indicating that the generated intent is generally well-regarded. Notably, math reasoning tasks achieve scores near 3 with agreement ratios approaching 100%, demonstrating that SWI-generated intent is particularly coherent, effective, and interpretable in mathematical problem-solving.
The relatively low scores observed in QA tasks may be attributed to the lack of multi-step guidance: the number of generated intents in QA tasks is often 1 or 2, while that in math and summarization is usually at least three. This finding indicates the advantages of multi-round iterative intents, although SWI can boost task performance even with few intents generated.

### v3

@@ 3.3 Baseline Settings
In § 5.2, we also investigate the synergy between SWI and existing LLM reasoning & planning methods, i.e., Chain-of-Thought (CoT) (Kojima et al., 2022) and Plan-and-Solve (PS) (Wang et al., 2023). CoT aims to elicit LLM reasoning using the answer-trigger prompt $\Phi_{i}^{\text{CoT}}$ as “Let’s think step by step”, while PS applies the following prompt $\Phi_{i}^{\text{PS}}$ to construct plans before problem-solving: “Let’s first understand the problem and devise a plan to solve the problem. Then, let’s carry out the plan and solve the problem step by step.” With such answer-trigger prompts $\Phi_{i}$, the generation process (Eq. 1) is given by
@@
4 Main Results
This section presents the experimental results to verify the effectiveness of SWI on diverse generation and reasoning tasks.
Method Text Summarization (Average ROUGE-1/2/L/Lsum %) Multi-task QA (Accuracy %) Math Reasoning (Accuracy %)
CDM XSum XL-Sum DialogSum WikiLingua BBH MMLU MMLU-Pro GSM8K GSM8K-P MATH500
w/o SWI 23.38 11.90 11.29 16.92 15.01 56.65 52.40 39.27 79.08 81.89 38.20
SWI 23.91 13.90 13.80 19.57 16.53 63.11 59.22 43.72 80.06 82.88 43.00
Table 3: Main results on text summarization, multi-task QA, and mathematical reasoning tasks.
@@ 4.1 Text Summarization
First, we demonstrate that SWI benefits natural language generation tasks like summarization by more explicitly analyzing the source document point by point and better planning the generation of the final summary.
@@
We evaluate the quality of summaries using the ROUGE score (Lin, 2004), which counts the overlaps of the generated summaries and reference summaries. Specifically, we average the ROUGE-1 (unigrams), ROUGE-2 (bigrams), ROUGE-L (longest common subsequences), and ROUGE-LSum (sentence-level ROUGE-L) scores as the final ROUGE score. As shown in Table 3, our SWI method consistently surpasses the direct generation baseline (“w/o SWI”), confirming its effectiveness in enhancing the quality of text summaries.
4.2 Multi-task Question Answering
Beyond text understanding and generation, we also consider multi-task QA tasks. We test the effect of SWI on three large-scale and challenging benchmarks designed as multiple-choice QA tasks, where the model is asked to select the most appropriate one from the given options to answer the question. Table 3 shows that our SWI method consistently improves the direct generation baseline by a large margin. The results demonstrate the efficacy of SWI in reasoning-intensive QA tasks.
4.3 Mathematical Reasoning
Additionally, we explore the efficacy of SWI on three high-quality math benchmarks. Unlike multiple-choice QA in § 4.2, where the model picks an option from the given list, math tasks require LLMs to generate numerical values as the answer. As shown in Table 3, SWI consistently improves the model performance over direct generation, showing its effectiveness in enhancing LLM on problem analysis and mathematical reasoning.
5 Analysis
Dataset Method Precision Recall F1
CDM w/o SWI 26.06 76.28 36.37
SWI 34.22 55.89 37.79
XSum w/o SWI 11.06 48.38 15.15
SWI 14.77 37.30 16.29
XL-Sum w/o SWI 8.96 61.88 13.79
SWI 12.96 46.72 16.51
DialogSum w/o SWI 23.99 57.08 29.55
SWI 34.92 45.19 31.20
WikiLingua w/o SWI 23.33 65.55 30.63
SWI 32.40 54.98 35.78
@@ 5 Analysis
Table 4: Fact Checking Evaluation of Summaries. We compare the atomic facts drawn from the LLM-generated summaries and the golden references, and compute recall, precision, and F1 scores (%).
5.1 Fact Checking of Summaries
LLMs frequently generate hallucinated content (Ji et al., 2023; Li et al., 2024a), which can not be detected by lexical metrics like ROUGE. To assess this issue, we adopt a more semantically sophisticated fact-checking metric (Hwang et al., 2025), which quantifies factual consistency by calibrating the extent of fabricated statements (low precision) and omitted factual information (low recall). Specifically, we use GPT-4o-mini (OpenAI, 2024) to decompose both generated and reference summaries into atomic fact sets and measure their coverage to quantify factual consistency.
We evaluate $100$ random samples from each summarization dataset using this fact-checking metric, with results presented in Table 4. Directly generated summaries (“w/o SWI”) tend to be more lengthy and verbose, resulting in higher recall scores. In contrast, SWI-generated summaries exhibit greater accuracy, conciseness, and factual correctness, with fewer hallucinations. Overall, SWI consistently outperforms the direct generation baseline in terms of F1 score.
Method GSM8K GSM8K-P MATH500 Avg.
CoT w/o SWI 77.86 80.07 42.00 66.64
SWI 80.21 82.55 42.80 68.52
PS w/o SWI 72.56 75.35 40.00 62.64
SWI 79.45 82.54 41.40 67.80
Table 5: LLM Reasoning & Planning with SWI. When additional LLM reasoning (CoT) and planning (PS) methods are adopted, the exact matching scores (%) on multiple math datasets with or without SWI.
Method News Article Dialogue Wiki Article
CDM XSum XL-Sum DialogSum WikiLingua
CoT w/o SWI 23.17 11.54 11.11 15.77 14.44
SWI 24.25 13.86 13.73 19.49 16.88
PS w/o SWI 24.12 12.21 11.91 17.92 15.86
SWI 24.43 12.46 12.28 18.95 16.76
Table 6: LLM Reasoning & Planning with SWI. When additional LLM reasoning (CoT) and planning (PS) methods are adopted, the ROUGE scores (%) on multiple summarization datasets with or without SWI.
5.2 Synergy with Other Methods
In recent years, various methods have been proposed to boost the reasoning and planning abilities of LLMs. Since our SWI method is orthogonal to previous work, it is necessary to compare the performance and study the synergy between SWI and them. As mentioned in § 3.3, we adopt representative LLM reasoning method Chain-of-Thought (CoT) (Kojima et al., 2022) and LLM planning method Plan-and-Solve (PS) (Wang et al., 2023).
When additional reasoning (CoT) and planning (PS) methods are adopted, Table 5 and Table 6 present the LLM performance on multiple math and summarization tasks, respectively. Comparing the results in Table 3 and Table 5, our SWI method (➁ in Table 3) beats CoT (➀ in Table 5) and PS (➂ in Table 5). Moreover, the combination of CoT+SWI (➁ in Tables 5 and 6) boosts the CoT method (i.e., ➁>>➀), and the synergy of PS+SWI (➃ in Tables 5 and 6) also improves the PS-alone performance (i.e., ➃>>➂). These results verify that SWI works synergistically with existing LLM reasoning & planning methods.
5.3 Generalizability to Different LLMs
To further validate the generalizability of our SWI method, we evaluate its effect on different sizes and types of LLMs. Aside from the results of LLaMA3-8B (Table 3), we present the results on multiple mathematical reasoning tasks using LLaMA3-3B (with 3B parameters) and LLaMA3-8B-R1, which is fine-tuned using reasoning data distilled from DeepSeek R1 (Guo et al., 2025). As observed in Table 7, our SWI method brings consistent improvements over the direct generation baseline, verifying the effectiveness of SWI when applied to models of different model sizes and LLM types (i.e., chat and reasoning models).
Model Method GSM8K GSM8K-P MATH500 Avg.
LLaMA3-3B w/o SWI 45.64 46.82 27.20 39.89
SWI 65.05 67.58 32.80 55.14
LLaMA3-8B-R1 w/o SWI 68.08 70.72 56.40 65.07
SWI 75.44 79.24 57.00 70.56
Table 7: Generalizability of SWI to different LLMs. When different sizes and types of LLMs are adopted, the exact matching scores (%) on multiple mathematical reasoning datasets with or without SWI.
Method News Article Dialogue Wiki Article
CDM XSum XL-Sum DialogSum WikiLingua
w/o SWI 23.38 11.90 11.29 16.92 15.01
SWI (V0) 23.91 13.90 13.80 19.57 16.53
SWI (V1) 24.27 14.12 14.10 19.43 17.34
SWI (V2) 24.04 14.69 14.66 19.24 17.06
SWI (V3) 24.17 13.83 14.02 18.82 16.10
Table 8: Results of Different SWI Prompt Variants. When different paraphrases of SWI prompts are adopted, the ROUGE scores (%) on various text summarization datasets with or without Speaking with Intent (SWI).
5.4 SWI Prompt Variants
In this work, we implement SWI in a straightforward prompting way for simplicity and reproducibility. To demonstrate that SWI works effectively irrespective of specific prompt design, we conduct experiments on different SWI prompt variants. Specifically, the original SWI prompt (“V0”) is paraphrased into three different versions (details in the appendix) by GPT-4o (OpenAI, 2024). Table 8 shows the performance of each SWI variants over multiple summarization tasks. As illustrated, the proposed SWI method, regardless of its prompt implementation, maintains a consistent advantage over the direct generation baseline, substantiating that SWI is an effective and general framework that brings intent—a key cognitive concept for reasoning and communication—into LLM generation.
Task Dataset # Input Tokens # Output Tokens
w/o SWI SWI $\Delta$ w/o SWI SWI $\Delta$
Sum CDM 920 1028 +108 166 434 +161%
XSum 542 650 +108 135 374 +177%
XL-Sum 613 721 +108 150 405 +170%
DialogSum 263 371 +108 77 238 +209%
WikiLingua 525 633 +108 145 386 +166%
QA BBH 203 313 +110 96 244 +154%
MMLU 193 303 +110 63 229 +263%
MMLU-Pro 270 380 +110 493 703 +43%
Math GSM8K 123 233 +110 211 280 +33%
GSM8K-P 123 233 +110 213 273 +28%
MATH500 133 243 +110 810 922 +14%
Table 9: Efficiency Study. The number of input and output tokens with or without SWI.
5.5 Efficiency Study
To study the efficiency cost of requiring LLMs to speak with intent, we present the number of input and output tokens with or without SWI in Table 9. SWI adds about $110$ extra input tokens in the system prompt (as in Table 1) for additional instructions, and the extra output tokens are not more than 300 tokens (percentages in Table 9) on average compared to direct generation. Hence, this cost should be traded off against the substantially improved performance, transparency, and interpretability of SWI.
@@ 5.5 Efficiency Study
Figure 3: Case Study. The source article, reference summary, and SWI output of a text summarization example.
5.6 Case Study
In addition to the math example in Figure 2, we conduct a case study on the summarization task to provide more insights into the benefits of SWI. Figure 3 showcases the SWI outputs, where the generated intent is well formulated and articulated, guiding a point-by-point summarization process that leads to a final summary that is accurate, concise, and abstractive, while effectively capturing the key information of the source article. This probably leads to the high Precision and F1 scores of SWI in Table 4, suggesting the validity of SWI in text generation tasks. Similarly, SWI enables LLM to have progressive planning when solving math problems. This is a critical ability when dealing with complex problems that require the divide-and-conquer strategy.
Task Dataset Total Unique Per Ins.
Sum CDM 53,086 86 4.6
XSum 50,697 92 4.5
XL-Sum 56,100 84 4.9
DialogSum 4,739 96 3.2
WikiLingua 12,463 91 4.2
QA BBH 7,425 49 1.3
MMLU 19,380 106 1.4
MMLU-Pro 21,916 122 1.8
Math GSM8K 5,237 45 4.0
GSM8K-P 4,719 43 3.9
MATH500 3,619 84 7.2
Table 10: Intent Statistics. The number of total and unique intents (verbs) of each dataset, as well as the average number of intents per instance (“Per Ins.”).
5.7 Intent Statistics
To further analyze the pattern and variability of the intents generated by our SWI method, we present the intent statistics across different tasks. Specifically, we extract and count the verbs in the generated intent statements, which are required to be in the format “To do something.”. Table 10 shows the number of total verbs (“Total”) and unique verbs (“Unique”) of each dataset, as well as the average number of intents per instance (“Per Ins.”). We observe that the number of unique intents in the summarization task is larger, indicating that summarizing documents demands a higher intent variability. In addition, summarization and math tasks generally have more intents per instance than QA, likely due to their longer outputs, as also observed in the efficiency study (Table 9). Among the three math datasets, MATH500 is relatively harder, as the model performance is lower in Table 3. Thus, solving the problems in MATH500 requires more thinking steps and longer reasoning chains, which is consistent with the observation that its number of unique intent verbs and the average number of intents per instance are larger than GSM8K.
Figure 4: The semantic distribution of intents across different task types and among all English verbs.
Furthermore, we investigate the intents across different task types and the distribution of intent verbs among all English verbs. First, we feed the same model that generates the outputs (i.e., LLaMA3-8B) with each verb and extract the last-layer hidden states, which indicate how the generator perceives and utilizes the intents. Then, t-SNE (van der Maaten and Hinton, 2008) is applied to visualize the semantic representations of each unique intent verb and all 11,531 English verbs drawn from WordNet (Miller, 1992). We observe from Figure 4(a) that all three tasks involve diverse intents, indicating the need for versatile intent skill sets when performing different tasks. Figure 4(b) illustrates that intent verbs mainly lie in a certain cluster, showing their specialty from common English verbs. In addition, Figure 5 presents the top 10 common intent verbs in summarization, QA, and math tasks, demonstrating the nuance of intents used in different tasks.
Figure 5: Top 10 common intent verbs of (a) summarization, (b) question answering, and (c) math tasks.
5.8 Intent Quality Evaluation
@@
Although we have shown that SWI boosts performance across a broad range of tasks, verifying the quality of generated intents is also significant. Thus, we hire human evaluators to assess the quality of generated intent across three criteria: coherence, effectiveness, and interpretability. Coherence measures how well the intent guides analysis and reasoning, effectiveness evaluates its contribution to problem-solving, and interpretability assesses its role in enhancing user understanding of the generated content. For each instance, human evaluators are provided with evaluation instructions, task input (e.g., the math problem, question with options, or source article), SWI-generated output, and assessment check boxes. They are then asked to evaluate the following aspects:
@@
  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for performing the task?
@@
Evaluation scores range from 1 (Bad), 2 (Fair), to 3 (Good). Agreement ratios are calculated as follows: 1 if all three evaluators agree, 0.5 if two agree, and 0 if all scores differ. As shown in Table 11, human evaluation scores for all aspects across datasets exceed 2.3, indicating that the generated intent is generally well-regarded. Notably, we observe very strong assessment scores (near 3) with substantial agreement (approaching 100%) for both summarization and math tasks, demonstrating that SWI-generated intent is particularly coherent, effective, and interpretable.
The relatively low (but still fairly good) scores observed in QA tasks may be attributed to the lack of multi-step guidance: as presented in Table 10, the average number of intents per instance in QA tasks is often 1 or 2, which is much lower than that in summarization and math tasks. This finding indicates the advantages of multi-round iterative intents, with SWI being able to boost task performance even with a few intents generated.
6 Related Work
Intent in NLG and LLMs.
Since the seminal work (Grosz and Sidner, 1986), intent has played a critical role in NLG. In the classical NLG pipeline (Reiter and Dale, 1997, 2000), content determination and document planning are modeled as a process of communicative goals decomposition and ordering, where the resulting planned communicative acts the NLG system wants to achieve are its intentions. Most approaches followed this framework by implementing a computational model of the Rhetorical Structure Theory (RST) (Mann and Thompson, 1988), a discourse theory that explains how parts of a multi-sentential text relate to each other functionally, i.e., how each piece serves a communicative purpose relative to the whole. For instance, RST has been applied in NLG to many genres, ranging from handling explanation dialogues (Moore and Paris, 1993) to generating persuasive evaluative arguments (Carenini and Moore, 2006). In all these applications, the text was planned before being generated and intentions were explicit.
In contrast, in modern LLM-driven generation, intent is typically implicit. In other words, what we see is only the generated text, with no access to the underlying communicative goals and corresponding intentions. In this respect, SWI can be seen as the first attempt to make LLM intentionality more explicit, bridging the gap between classic NLG and LLMs, arguably boosting LLMs’ controllability and interpretability.
LLM Reasoning.
Early LLMs were rather poor at reasoning (Radford et al., 2018, 2019), and scaling pre-training was shown not to be a feasible solution for improving reasoning (Chu et al., 2025). Instead, Chain-of-Thought (CoT) prompting (Kojima et al., 2022; Wei et al., 2022) demonstrated that by modifying the prompt, LLMs can elicit a beneficial step-by-step reasoning process at test time without additional training (Li et al., 2024b; Yeo et al., 2025; Zhang et al., 2025). Building on CoT, various reasoning techniques have emerged (Xu et al., 2025). Among them, a recent research ARR (Yin and Carenini, 2025) consistently outperforms CoT on multiple QA tasks, where analyzing the intent of questions is its most effective component. Different from ARR, we enable LLMs to articulate their intent, using it to guide subsequent analysis and reasoning for improved task performance.
Inspired by the success of CoT and similar prompting techniques, very recent research is increasingly focusing on enhancing LLMs reasoning abilities by explicitly training them for reasoning using reinforcement learning (RL) (Sutton and Barto, 2018; Shao et al., 2024). Intriguingly, the success of SWI demonstrated in this paper could spur a similar explosion of research on training LLMs with RL (Wang et al., 2025; Setlur et al., 2025) to better analyze and formulate intentions.
Intent-related Research.
Intent Detection (ID) and New Intent Discovery (NID) (Kumar et al., 2022; Liang et al., 2024; Zhang et al., 2024a; Tang et al., 2024; Zhang et al., 2024c; Qian et al., 2024; Yin et al., 2025), which classify utterances into known or novel intent categories, are longstanding challenges in natural language understanding (Larson et al., 2019; Casanueva et al., 2020; Zhang et al., 2021; Weld et al., 2022). Typically, these tasks are approached as classification problems (Wang et al., 2024a; Yoon et al., 2024; Zhang et al., 2024b; Sakurai and Miyao, 2024), where models assign sentences to predefined intent classes. On the contrary, our SWI method generates intent as free-form text rather than fixed categories, enhancing flexibility and fluency. SWI naturally integrates intent statements as planning into the reasoning process, providing contextual guidance for subsequent analysis.